Explain Answer

Answer all Parts

Assume a major investment service has just given Oasis Electronics its highest investment rating, along with a strong buy recommendation. As a result, you decide to take a ook for yourself and to place a value on the companys stock. Heres what you find: This year, Oasis paid its stockholders an annual dividend of $3.00 a share, but because of its high rate of growth in earnings, its dividends are expected to grow at the rate of 12% a year for the next 4 years and then to level out at 9% a year. So far, youve learned that the stock has a beta of 1.80, the nsk-free rate of return is 5%, and the expected return on the market is 1 1% Using the CAPM to find the required rate of return, put a value on this stock. Using the CAPM, the required rate of return on the investment is 96 Round to two decimal places. The value of the companys stock is $(Round to the nearest cent.)
Answer #1

Solution:

Required rate of return=Risk free rate+Beta*(Market rate-Riskfree rate)

=5+1.80*(11-5)=15.80%

Dividend for year 1 = (3*1.12) = $3.360

Dividend for year 2 = (3.360*1.12) = $3.7632

Dividend for year 3= (3.7632*1.12) = $4.214784

Dividend for year 4= (4.214784*1.12) = $4.72055808

Value after year 4= (Dividend for year 4*Growth rate)/(Requiredreturn-Growth rate)

=(4.72055808*1.09)/(0.1580-0.09)

=$75.667769222

Current price=Future dividends*Present value of discountingfactor(15.80%,time period)

= 3.360/1.1580+3.7632/1.1580^2+4.214784/1.1580^3+4.72055808/1.1580^4+75.667769222/1.1580^4

which is equal to

$53.13 (Approx).